A car running at a speed of 72 km/h is to be stopped in exactly 10 seconds. Find the uniform deceleration needed.

1
Expert's answer
2021-12-27T08:01:11-0500

Given quantities:

"v_o = 72 \frac{km}{h}=20 \frac{m}{s}"

"v = 0"

"t = 10s"

"a - ?"

"a = \large\frac{v-v_o}{t}=-2\frac{m}{s^2}"

the uniform deceleration is "-2\frac{m}{s^2}"